# BatteryLife: A Comprehensive Dataset and Benchmark for Battery Life Prediction

### v6

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737372isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Data mining
@@ 1. Introduction
CALCE (Xing et al., 2013) 13 1 1 1 2 Lab test Li-ion 2013
@@
SNL (Preger et al., 2020) 52 1 3 3 23 Lab test Li-ion 2020
MICH_EXP (Mohtat et al., 2021) 12 1 1 3 6 Lab test Li-ion 2021
@@
BatteryML (Zhang et al., 2023a) 383 2 5 5 186 Lab test Li-ion 2024
@@
  • Limited dataset size. As shown in Table 1, the previous largest dataset is BatteryML (Zhang et al., 2023a) that offers 383 end-of-life batteries by integrating 7 public datasets (Attia et al., 2020; Severson et al., 2019; Ma et al., 2022; Devie et al., 2018; Preger et al., 2020; Juarez-Robles et al., 2020; Li et al., 2021; Juarez-Robles et al., 2021). However, this undertaking overlooks several datasets (Zhu et al., 2022; Weng et al., 2021; Mohtat et al., 2021) and lacks datasets (Cui et al., 2024; Wang et al., 2024c; Li et al., 2024) that are released recently. In this vein, BatteryML, whose size is smaller than half of the sum of existing datasets, fails to provide comprehensive insights into modern battery life data.
@@
To address these issues, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. From a dataset perspective, BatteryLife stands out as the largest and most diverse battery life dataset by integrating 16 datasets. Specifically, BatteryLife is 2.5 times the size of BatteryML (Zhang et al., 2023a) in terms of battery number, covering 99,000 samples. Remarkably, BatteryLife furnishes unparalleled diversity, delivering 4 times format, 11.8 times chemical system, 1.8 times operation temperature, and 2.2 times charge/discharge protocol compared to BatteryML (Zhang et al., 2023a). Notably, BatteryLife pioneers as the first work that releases zinc-ion (Zn-ion) and sodium-ion (Na-ion) battery life datasets, and additionally offers large-capacity Li-ion batteries that are tested by the battery manufacturer (i.e., CALB Group Co., Ltd.). The formidable diversity challenges the model’s capability to learn essential mappings from early data to life and allows inspiration for the data mining community and a broader battery community.
Figure 1. Partial cycling data (voltage and current time series) from three batteries with the same specifications tested under a protocol.
[…]
[zero figure]Partial cycling data (voltage and current time series) from three batteries with the same specifications tested under a protocol.
@@
  • The largest battery life dataset: BatteryLife offers 99,000 samples from 990 batteries, which is 2.5 times the size of the previous largest one. The unprecedented dataset size allows researchers to generate practical insights into battery life data by making advances on BatteryLife.
@@ 2.1. Degradation Test
Figure 2. Partial cycling data in degradation tests from (a) a lab-tested Li-ion battery, (b) a lab-tested Zn-ion battery, (c) a lab-tested Na-ion battery, and (d) an industry-tested Li-ion battery. The positive and negative values in current represent charging and discharging, respectively. The gray dotted lines are separators of different cycles. The charging and discharging periods are labeled in the first cycle as illustrations.
[…]
[first figure]Partial cycling data in degradation tests from (a) a lab-tested Li-ion battery, (b) a lab-tested Zn-ion battery, (c) a lab-tested Na-ion battery, and (d) an industry-tested Li-ion battery. The positive and negative values in current represent charging and discharging, respectively. The gray dotted lines are separators of different cycles. The charging and discharging periods are labeled in the first cycle as illustrations.
@@ 3.1. Dataset Overview
Figure 3. SOH degradation trajectories in BatteryLife.
[…]
[SOH trajectories]SOH degradation trajectories in BatteryLife.
@@ 3.2. Popular Benchmark Methods
Figure 4. Overview of the CyclePatch framework.
[…]
[Cyclepatch]Overview of the CyclePatch framework.
@@ 3.3. CyclePatch
(3) $\left[\textbf{X}_{1},\textbf{X}_{2},\textbf{X}_{3},\cdots,\textbf{X}_{S}\right]={\rm Segment}(\textbf{X}_{i:S}),$
@@ 4. Experiments
Figure 5. The performance of the top-three models with an increasing number of usable cycles.
[…]
[error curve]The performance of the top-three models with an increasing number of usable cycles.
@@ 4.1. Experimental Setup
We randomly split BatteryLife into training, validation, and testing sets with a 6:2:2 ratio. All models are trained to minimize mean squared error using the Adam optimizer (Kingma and Ba, 2017). Note that models utilizing CyclePatch are prefixed with ”CP” in their names. For instance, CPTransformer employs CyclePatch and uses a Transformer encoder as the inter-cycle encoder. Transformer refers to the Transformer encoder in this work. We conduct more than 3,000 hyperparameter searches in total (details in Appendix B.1), and select the best-performing hyperparameters based on the validation set. To reduce the impact of experimental fluctuations, we run each experiment three times and report the mean±$\pm$±standard deviation on testing sets for all evaluations.
@@ 4.4. Ablation Studies (RQ2)
We conduct ablation studies of the proposed CyclePatch to study the effectiveness of its key components, and present the results in Table 4. Specifically, we study CPMLP and CPTransformer considering their leading performance across datasets. ”w/o Intra” refers to the removal of the intra-cycle encoder; ”w/o Inter” indicates the absence of the inter-cycle encoder; MLP and Transformer serve as naive baselines without patching. From Table 4, it is evident that all components in CyclePatch significantly contribute to its performance. Concretely, the incorporation of the intra-cycle encoder enhances the model’s ability to model interactions among variables within cycles, leading to performance gains across all datasets. Removing the inter-cycle encoder reduces performance, highlighting the importance of modeling inter-cycle interactions. Moreover, it is observed that CPMLP without either intra-encoder or inter-encoder performs worse than MLP on Na-ion and CALB datasets, whereas MLP naively models the intra-cycle and inter-cycle information. This indicates that modeling intra-cycle and inter-cycle information are both critical. These results collectively validate the effectiveness of the CyclePatch design.
@@ 4.6. Cross-domain Transferability (RQ4)
In real-world situations, new battery types are often outnumbered by older ones, and Li-ion batteries undergo continuous industrial improvements. Therefore, developing a model with cross-domain transferability is crucial. In this work, we evaluate the cross-domain transferability of models by investigating if a model pretrained on Li-ion dataset can be transferred to other domains, and report the results in Table 6. Specifically, we study the transferability of CPMLP which is the best model in Li-ion domain. Three transferability evaluations are considered: (1) Frozen: Directly applying the pretrained CPMLP on target domains. (2) Fine-tuning (FT): Fine-tune all parameters of the pretrained CPMLP. (3) Domain adaptation (DA): we implement a domain adaptation technique by minimizing maximum mean discrepancy (Han et al., 2022) between the target-domain and source-domain embeddings produced by the last hidden layer. By expectation, with transfer learning techniques, an effective transferrable model should outperform the best baseline that is trained using only target-domain training data.
@@ 6. Acknowledgments
Appendix A Further Details of Datasets
@@ A.1. Data Preprocessing
In this section, we introduce the preprocessing details for all datasets. We obtained HNEI, MICH, MICH_EXP, SNL, and UL_PUR datasets from BatteryArchive. Other previously available datasets were sourced from the links shared in the original papers.
For all previously available datasets, the present study filtered out the batteries whose SOH is not degraded to $\lambda+2.5\%$ ($\lambda=90\%$ for the CALB dataset and $\lambda=80\%$ for other datasets), and then conducted the data statistics. Note that (1) BatteryArchive is an active website, and the data statistics were conducted on the available data until September 2024. (2) Stanford (Cui et al., 2024) were released multiple times and the data used in this work is from the first release. As for BatteryLife, we further excluded the batteries whose life labels are no larger than 100. Additionally, we found that some batteries in the UL_PUR dataset contain sudden significant SOH drops without recovery. We conjecture that these drops are caused by equipment faults and thus removed these batteries. The data statistics of BatteryLife were conducted on the remaining batteries.
Furthermore, some outliers might exist in the cycling data. For the 7 datasets standardized by BatteryML (Zhang et al., 2023a), we detected abnormal cycles by running a median filtering method on discharge capacity degradation trajectory following BatteryML (Zhang et al., 2023a) and removed the detected outliers. As for other datasets, we first removed the rest performance tests (RPT) and formation cycles. The RPT cycles are removed because they comply with charge/discharge protocols that are different from the cycling protocols, and the RPT cycles are generally detected as outliers by BatteryML (Zhang et al., 2023a). Likewise, formation cycles are removed. Moreover, some cycles of sudden rise or drop in discharge capacities are manually removed. After data cleaning, each battery was stored in a data format consistent with BatteryML (Zhang et al., 2023a).
@@ A.2. Degradation Test Details
The data for the zinc-ion batteries were collected using custom-made coin batteries fabricated in the laboratory. The coin batteries utilized stainless steel CR2032, CR2025 and CR2016 battery cases sourced from Guangdong Canrd New Energy Technology Co., Ltd., paired with stainless steel spring washers with a height of 0.6 mm and stainless steel gaskets of 0.5 mm thickness, all sealed under a pressure of 75 kg/cm². The cathode material of the zinc battery was commercial manganese dioxide cathodes purchased from Guangdong Canrd New Energy Technology Co., Ltd. The anode material consisted of 99.99% pure zinc foil produced by a local manufacturer. The separator used was Glass Fiber-D. The electrolyte employed was an aqueous solution of 2 mol/L zinc sulfate and 0.1 mol/L manganese sulfate, with various additives of different types and concentrations. The battery tests were conducted at a current density of 400 mA/m² and evaluated at temperatures of 25°C, 30°C, and 40°C. The discharge capacity of the 10th cycle is used as nominal capacity.
The high-rate sodium-ion 18650 cylindrical batteries utilized in this study were sourced from Zhuhai Punashidai New Energy Co., Ltd., and were employed without any additional processing. The specific composition of the electrode materials and electrolyte remains proprietary information of the manufacturer. The nominal capacity (1.0 Ah) of these batteries was determined through charge and discharge cycles at a 1C rate. All cycling tests were conducted within a voltage range of 2.0 V to 4.0 V for both charging and discharging processes. To assess the long-term performance, cycling tests were carried out at 12 distinct current rates, ranging from 2C to 6C, under one operation temperature: 25°C.
@@ A.3. Overview of The Integrated Datasets
  • CALCE (Xing et al., 2013): This dataset consists of 13 prismatic lithium-ion batteries, with $\rm LiCoO_{2}$ as the positive electrode and graphite as the negative electrode, while the electrolyte is unknown. The nominal capacity is 1.1 Ah. The working temperature of all batteries is 25 degrees Celsius. 2 different protocols are included totally in this dataset.
@@
  • HNEI (Devie et al., 2018): In this dataset, there are 14 18650-format lithium-ion batteries (LG Chem, ICR18650 C2) whose cathode is a mixture of $\rm LiCoO_{2}$ and $\rm LiNi_{0.4}Co_{0.4}Mn_{0.2}O_{2}$, and anode is graphite. The electrolyte composition is not disclosed, and nominal capacity is 2.8 Ah for all batteries. There are 2 different protocols with a working temperature of 25 degrees Celsius.
@@
  • MATR (Attia et al., 2020; Severson et al., 2019): This dataset consists of 169 commercial Li-ion batteries in a format of 18650 cylindrical batteries whose positive electrode is lithium iron phosphate ($\rm LiFePO_{4}$) and negative electrode is graphite. The the electrolyte is unknown. The nominal capacity of all batteries is 1.1 Ah. All batteries were cycled under a working temperature of 30 degrees Celsius with varied fast-charging protocols and identical discharge protocols. Totally, 81 different fast-charging protocols are included in this dataset.
@@
  • UL_PUR (Juarez-Robles et al., 2020): This dataset covers 10 batteries in a format of 18650 cylindrical batteries. The cathode is $\rm LiNi_{0.8}Co_{0.15}Al_{0.05}O_{2}$. The anode is graphite. The electrolyte formula is unknown. The nominal capacity is 3.4 Ah. The working temperature is 23 degrees Celsius. All batteries underwent the same cycling protocol.
@@
  • SNL (Preger et al., 2020): This dataset consists of 52 batteries in the format of 18650 cylindrical batteries. The positive electrode covers $\rm LiFePO_{4}$, $\rm LiNi_{0.81}Co_{0.14}Al_{0.05}O_{2}$, and $\rm LiNi_{0.84}Mn_{0.06}Co_{0.1}O_{2}$, while the negative electrode is graphite. The electrolyte formula is unknown. There are 3 nominal capacities consisting of 1.1 Ah, 3.2 Ah, and 3.0 Ah. There are 23 different protocols in this dataset with 3 different working temperatures (15°C, 25°C and 35°C).
@@
  • RWTH (Li et al., 2021): This dataset has 48 lithium-ion batteries in the format of 18650 cylindrical batteries. The positive electrode is lithium nickel manganese cobalt oxide (NMC) and the negative electrode is carbon. The electrolyte is unknown. The nominal capacity is 3 Ah. All batteries are cycled with the same protocols with a working temperature of 25 degrees Celsius.
@@
  • MICH_EXP (Mohtat et al., 2021): This dataset consists of 12 lithium-ion batteries in the format of pouch batteries. The positive electrode is $\rm LiNi_{1/3}Co_{1/3}Mn_{1/3}O_{2}$ (NCM111), and the negative electrode is $\rm graphite$. The electrolyte consists of 1M $\rm LiPF_{6}$ salt in solvents of ethylene carbonate (EC) and ethyl methyl carbonate (EMC) with a ratio of 3:7. The nominal capacity is 5.0 Ah. There are 6 different protocols with 3 different working temperatures in total (-5°C, 25°C, and 45°C).
[…]
@@
  • HUST (Ma et al., 2022): This dataset consists of 77 lithium-ion batteries in the format of 18650 cylindrical batteries. The positive electrode is lithium-iron-phosphate, and the negative electrode is graphite. The electrolyte for this dataset is unknown. The nominal capacity is 1.1 Ah. There are 77 different multi-stage discharge protocols but an identical fast-charging protocol with 30 degrees Celsius as the working temperature.
@@
  • Stanford (Cui et al., 2024): This dataset covers 41 batteries in the format of pouch batteries whose positive electrode is $\rm LiNi_{0.5}Mn_{0.3}Co_{0.2}O_{2}$, negative electrode is artificial graphite. The electrolyte is a mixture of $\rm LiPF_{6}$, ethylene carbonate, ethyl methyl carbonate, dimethyl carbonate, and vinylene carbonate. The nominal capacity of these batteries is 0.24 Ah. All batteries underwent different formation protocols but were cycled with an identical charge/discharge protocol under a working temperature of 30 degrees Celsius.
  • XJTU (Wang et al., 2024c): There are 23 batteries in this dataset in the format of 18650 cylindrical batteries. For all batteries, the positive electrode is $\rm LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}$ and the negative electrode is graphite. The electrolyte is unknown. The nominal capacity is 2 Ah. This dataset includes 2 different operation conditions with a working temperature of 20 degrees Celsius.
  • ISU_ILCC (Li et al., 2024): This dataset comprises 240 Li-polymer batteries in the format of 502030-size pouch batteries. The positive electrode is lithium nickel manganese cobalt oxide, and the negative electrode is graphite. The electrolyte remains unknown. And the nominal capacity is 0.25 Ah. There are 202 different charge/discharge protocols in this dataset, with 30 degrees Celsius as the working temperature.
  • Zn-ion: This dataset consists of 95 zinc-ion batteries in the format of coin batteries. The positive electrode is $\rm MnO_{2}$ while the negative electrode is zinc metal. The electrolyte is unknown. The discharge capacity of the $10^{th}$ cycle was taken as the nominal capacity for each battery. All batteries underwent the same charge/discharge protocol at 3 different working temperatures.
  • Na-ion: This dataset consists of 31 sodium-ion batteries in the format of 18650 cylindrical batteries. The positive electrode and the negative electrode are not disclosed. The electrolyte remains unknown. And the nominal capacity is 1.0 Ah. There are 12 different charge/discharge protocols in this dataset at 25 degrees Celsius.

### v7

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; August 03–07, 2025; Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data mining
@@ 1. INTRODUCTION
CALCE (Xing et al., 2013; He et al., 2011) 13 1 1 1 2 Lab test Li-ion 2013
@@
SNL (Preger et al., 2020) 50 1 3 3 23 Lab test Li-ion 2020
MICH_EXP (Mohtat et al., 2021) 18 1 1 3 6 Lab test Li-ion 2021
@@
BatteryML (Zhang et al., 2023a) 381 2 5 5 186 Lab test Li-ion 2024
@@
  • Limited dataset size. As shown in Table 1, the previous largest dataset is BatteryML (Zhang et al., 2023a) that offers 381 end-of-life batteries by integrating 7 public datasets (Attia et al., 2020; Severson et al., 2019; Ma et al., 2022; Devie et al., 2018; Preger et al., 2020; Juarez-Robles et al., 2020; Li et al., 2021; Juarez-Robles et al., 2021). However, this undertaking overlooks several datasets (Zhu et al., 2022; Weng et al., 2021; Mohtat et al., 2021) and lacks datasets (Cui et al., 2024; Wang et al., 2024c; Li et al., 2024) that are released recently. In this vein, BatteryML, whose size is smaller than half of the sum of existing datasets, fails to provide comprehensive insights into modern battery life data.
@@
To address these issues, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. From a dataset perspective, BatteryLife stands out as the largest and most diverse battery life dataset by integrating 16 datasets. Specifically, BatteryLife is 2.5 times the size of BatteryML (Zhang et al., 2023a) in terms of battery number, covering 90,000 samples. Remarkably, BatteryLife furnishes unparalleled diversity, delivering 4 times format, 11.8 times chemical system, 1.8 times operation temperature, and 2.2 times charge/discharge protocol compared to BatteryML (Zhang et al., 2023a). Notably, BatteryLife pioneers as the first work that releases zinc-ion (Zn-ion) and sodium-ion (Na-ion) battery life datasets, and additionally offers large-capacity Li-ion batteries that are tested by the battery manufacturer (i.e., CALB Group Co., Ltd.). The formidable diversity challenges the model’s capability to learn essential mappings from early data to life and allows inspiration for the data mining community and a broader battery community.
Figure 1. Partial cycling data (voltage and current time series) from three batteries with the same specifications tested under a protocol.: zero figure
@@
  • The largest battery life dataset: BatteryLife offers 90,000 samples from 990 batteries, which is 2.5 times the size of the previous largest one. The unprecedented dataset size allows researchers to generate practical insights into battery life data by making advances on BatteryLife.
@@ 2.1. Degradation Test
Figure 2. Partial cycling data in degradation tests from (a) a lab-tested Li-ion battery, (b) a lab-tested Zn-ion battery, (c) a lab-tested Na-ion battery, and (d) an industry-tested Li-ion battery. The positive and negative values in current represent charging and discharging, respectively. The gray dotted lines are separators of different cycles. The charging and discharging periods are labeled in the first cycle as illustrations.: first figure
@@ 3.1. Dataset Overview
Figure 3. SOH degradation trajectories in BatteryLife.: SOH trajectories
@@ 3.2. Popular Benchmark Methods
Figure 4. Overview of the CyclePatch framework.: Cyclepatch
@@ 3.3. CyclePatch
(3) $\left[\textbf{X}_{1},\textbf{X}_{2},\textbf{X}_{3},\cdots,\textbf{X}_{S}\right]={\rm Segment}(\textbf{X}_{1:S}),$
@@ 4. EXPERIMENTS
Figure 5. The performance of the top-three models with an increasing number of usable cycles.: error curve
@@ 4.1. Experimental Setup
[…]
@@ 4.4. Ablation Studies (RQ2)
We conduct ablation studies of the proposed CyclePatch to study the effectiveness of its key components, and present the results in Table 4. Specifically, we study CPMLP and CPTransformer considering their leading performance across datasets. ”w/o Intra” refers to the removal of the intra-cycle encoder; ”w/o Inter” indicates the absence of the inter-cycle encoder; MLP and Transformer severs as naive baselines without patching. From Table 4, it is evident that all components in CyclePatch significantly contribute to its performance. Concretely, the incorporation of the intra-cycle encoder enhances the model’s ability to model interactions among variables within cycles, leading to performance gains across all datasets. Removing the inter-cycle encoder reduces performance, highlighting the importance of modeling inter-cycle interactions. Moreover, it is observed that CPMLP without either intra-encoder or inter-encoder performs worse than MLP on Na-ion and CALB datasets, whereas MLP naively models the intra-cycle and inter-cycle information. This indicates that modeling intra-cycle and inter-cycle information are both critical. These results collectively validate the effectiveness of the CyclePatch design.
@@ 4.6. Cross-domain Transferability (RQ4)
In real-world situations, batteries of new types are often fewer than batteries of old types so that it is important to develop a model that has cross-domain transferability against the domain shifts caused by battery types. In this work, we evaluate the cross-domain transferability of models by investigating if a model pretrained on Li-ion dataset can be transferred to other domains, and report the results in Table 6. Specifically, we study the transferability of CPMLP which is the best model in Li-ion domain. Three transferability evaluations are considered: (1) Frozen: Directly applying the pretrained CPMLP on target domains. (2) Fine-tuning (FT): Fine-tune all parameters of the pretrained CPMLP. (3) Domain adaptation (DA): we implement a domain adaptation technique by minimizing maximum mean discrepancy (Han et al., 2022) between the target-domain and source-domain embeddings produced by the last hidden layer. By expectation, with transfer learning techniques, an effective transferrable model should outperform the best baseline that is trained using only target-domain training data.
@@ 6. ACKNOWLEDGMENTS
Appendix A FURTHER DETAILS ABOUT THE DATASETS
@@ A.1. Data Preprocessing
In this section, we introduce the preprocessing details for all datasets. We obtained HNEI, MICH, MICH_EXP, SNL, and UL_PUR datasets from BatteryArchive. Other datasets were sourced from the links shared in the original papers. For all datasets, the present study filtered out the batteries whose SOH is not degraded to $\lambda+2.5\%$ ($\lambda=90\%$ for the CALB dataset and $\lambda=80\%$ for other datasets), and then conducted the data statistics. Note that (1) BatteryArchive is an active website, and the data statistics were conducted on the available data until September 2024 in this paper. (2) Stanford (Cui et al., 2024) were released multiple times and the data used in this work is from the first release. As for BatteryLife, we further excluded the batteries whose life labels are no larger than 100. Additionally, we found that some batteries in the UL_PUR dataset contain sudden significant SOH drops without recovery. We conjecture that these drops are caused by equipment faults and thus removed these batteries. The data statistics were conducted on the remaining batteries. The statistics of BatteryLife were conducted on the remaining batteries.
Furthermore, some outliers might exist in the cycling data. For the 7 datasets standardized by BatteryML (Zhang et al., 2023a), we detected abnormal cycles by running a median filtering method on discharge capacity degradation trajectory following BatteryML (Zhang et al., 2023a) and removed the detected outliers. As for other datasets, we first removed the rest performance tests (RPT) and formation cycles. The RPT cycles are removed because they comply with charge/discharge protocols that are different from the cycling protocols, and the RPT cycles are also generally detected as outliers by BatteryML (Zhang et al., 2023a). Likewise, formation cycles are removed. Moreover, some cycles of sudden rise or drop in discharge capacities are manually removed. After data cleaning, each battery was stored in a data format consistent with BatteryML (Zhang et al., 2023a).
@@ A.2. Degradation Test Details
The data for the zinc-ion batteries were collected using custom-made coin batteries fabricated in the laboratory. The coin batteries utilized stainless steel CR2032, CR2025 and CR2016 battery cases sourced from Guangdong Canrd New Energy Technology Co., Ltd., paired with stainless steel spring washers with a height of 0.6 mm and stainless steel gaskets of 0.5 mm thickness, all sealed under a pressure of 75 kg/cm². The cathode material of the zinc battery was commercial manganese dioxide cathodes purchased from Guangdong Canrd New Energy Technology Co., Ltd. The anode material consisted of 99.99% pure zinc foil produced by a local manufacturer. The separator used was Glass Fiber-D. The electrolyte employed was an aqueous solution of 2 mol/L zinc sulfate and 0.1 mol/L manganese sulfate, with various additives of different types and concentrations. The battery tests were conducted at a current density of 400 mA/m² and evaluated at temperatures of 25°C, 30°C, and 40°C.
The high-rate sodium-ion 18650 cylindrical batteries utilized in this study were sourced from Zhuhai Punashidai New Energy Co., Ltd., and were employed without any additional processing. The specific composition of the electrode materials and electrolyte remains proprietary information of the manufacturer. The nominal capacity of these batteries was determined through charge and discharge cycles at a 1C rate. All cycling tests were conducted within a voltage range of 2.0 V to 4.0 V for both charging and discharging processes. To assess the long-term performance, cycling tests were carried out at 12 distinct current rates, ranging from 2C to 6C, under one operation temperature: 25°C.
@@ A.3. Overview of The Integrated Datasets
  • CALCE (Xing et al., 2013; He et al., 2011): This dataset consists of 13 prismatic lithium-ion batteries, with $\rm LiCoO_{2}$ as the positive electrode and graphite as the negative electrode, while the electrolyte is unknown. The nominal capacity is 1.1 Ah or 1.35 Ah. The working temperature of all batteries is 25 degree Celsius. 2 different protocols are included totally in this dataset.
@@
  • HNEI (Devie et al., 2018): In this dataset, there are 14 18650-format lithium-ion batteries (LG Chem, ICR18650 C2) whose cathode is a mixture of $\rm LiCoO_{2}$ and $\rm LiNi_{0.4}Co_{0.4}Mn_{0.2}O_{2}$, and the anode is graphite. The electrolyte composition is not disclosed, and the nominal capacity is 2.8 Ah for all batteries. There are 2 different protocols with a working temperature of 25 degree Celsius.
@@
  • MATR (Attia et al., 2020; Severson et al., 2019): This dataset consists of 169 commercial Li-ion batteries in a format of 18650 cylindrical batteries whose positive electrode is lithium iron phosphate ($\rm LiFePO_{4}$), the negative electrode is graphite, and the electrolyte is unknown. The nominal capacity of all batteries is 1.1 Ah. All batteries were cycled under a working temperature of 30 degree Celsius with varied fast-charging protocols and identical discharge protocols. Totally, 81 different fast-charging protocols are included in this dataset.
@@
  • UL_PUR (Juarez-Robles et al., 2020): This dataset covers 10 batteries in a format of 18650 cylindrical batteries. The cathode is $\rm LiNi_{0.8}Co_{0.15}Al_{0.05}O_{2}$. The anode is graphite. The electrolyte formula is unknown. The nominal capacity is 3.4 Ah. The working temperature is 23 degree Celsius. All batteries underwent the same cycling protocol.
@@
  • SNL (Preger et al., 2020): This dataset consists of 50 batteries in the format of 18650 cylindrical batteries. The positive electrode covers $\rm LiFePO_{4}$, $\rm LiNi_{0.81}Co_{0.14}Al_{0.05}O_{2}$, and $\rm LiNi_{0.84}Mn_{0.06}Co_{0.1}O_{2}$, while the negative electrode is graphite. The electrolyte formula is unknown. There are 3 nominal capacities consisting of 1.1 Ah, 3.2 Ah, and 3.0 Ah. There are 23 different protocols in this dataset with 3 different working temperatures (15°C, 25°C and 35°C).
@@
  • RWTH (Li et al., 2021): This dataset has 48 lithium-ion batteries in the format of 18650 cylindrical batteries. The positive electrode is lithium nickel manganese cobalt oxide (NMC), and the negative electrode is carbon. The electrolyte is unknown. The nominal capacity is 3 Ah. All batteries are cycled with the same protocols with a working temperature of 25 degree Celsius.
@@
  • MICH_EXP (Mohtat et al., 2021): This dataset consists of 18 lithium-ion batteries in the format of pouch batteries. The positive electrode is $\rm LiNi_{1/3}Co_{1/3}Mn_{1/3}O_{2}$ (NCM111), and the negative electrode is $\rm graphite$. The electrolyte consists of 1M $\rm LiPF_{6}$ salt in solvents of ethylene carbonate (EC) and ethyl methyl carbonate (EMC) with a ratio of 3:7. The nominal capacity is 5.0 Ah. There are 6 different protocols with 3 different working temperatures in total (-5°C, 25°C, and 45°C).
@@
  • MICH (Weng et al., 2021): This dataset consists of 40 lithium-ion batteries in the format of pouch batteries. The positive electrode is $\rm NCM111$ and the negative electrode is $\rm graphite$. The electrolyte consists of 1.0 M $\rm LiPF_{6}$ salt in solvents of EC and EMC with a ratio of 3:7 together with 2 wt% additive of vinylene carbonate. The nominal capacity is 2.36 Ah. There are 2 different protocols with 2 different working temperatures (25°C and 45°C).
  • HUST (Ma et al., 2022): This dataset consists of 77 lithium-ion batteries in the format of 18650 cylindrical batteries. The positive electrode is lithium-iron-phosphate, and the negative electrode is graphite. The electrolyte for this dataset is unknown. The nominal capacity is 1.1 Ah. There are 77 different multi-stage discharge protocols but an identical fast-charging protocol with 30 degree Celsius as the working temperature.
@@
  • Stanford (Cui et al., 2024): This dataset covers 41 batteries in the format of pouch batteries whose positive electrode is $\rm LiNi_{0.5}Mn_{0.3}Co_{0.2}O_{2}$, negative electrode is artificial graphite. The electrolyte is a mixture of $\rm LiPF_{6}$, ethylene carbonate, ethyl methyl carbonate, dimethyl carbonate, and vinylene carbonate. The nominal capacity of these batteries is 0.24 Ah. All batteries underwent different formation protocols but were cycled with an identical charge/discharge protocol under a working temperature of 30 degree Celsius.
  • XJTU (Wang et al., 2024c): There are 23 batteries in this dataset in the format of 18650 cylindrical batteries. For all batteries, the positive electrode is $\rm LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}$ and the negative electrode is graphite. The electrolyte is unknown. The nominal capacity is 2 Ah. This dataset includes 2 different operation conditions with a working temperature of 20 degree Celsius.
  • ISU_ILCC (Li et al., 2024): This dataset comprises 240 502030-size Li-polymer batteries. The positive electrode is lithium nickel manganese cobalt oxide, and the negative electrode is graphite. The electrolyte remains unknown. And the nominal capacity is 0.25 Ah. There are 202 different charge/discharge protocols in this dataset, with 30 degree Celsius as the working temperature.
  • Zn-ion: This dataset consists of 100 zinc-ion batteries in the format of coin batteries. The positive electrode is $\rm MnO_{2}$ while the negative electrode is zinc metal. The electrolyte is unknown. The discharge capacity of the $10^{th}$ cycle was taken as the nominal capacity for each battery. All batteries underwent the same charge/discharge protocol at 3 different working temperatures.
  • Na-ion: This dataset consists of 31 sodium-ion batteries in the format of 18650 cylindrical batteries. The positive electrode and the negative electrode are not disclosed. The electrolyte remains unknown. And the nominal capacity is 1.0 Ah. There are 12 different charge/discharge protocols in this dataset at 25 degree Celsius.